sue and david are siblings their combined age is 22 sue is four years old than david how old are sue and david
Answer by sophxmai(62)

Sue and David are siblings, their combined age is 22. Sue is four years older than David, how old are Sue and David.

If Sue and David's ages combined equal 22,
then let 22=x+y. (x is David's age, y is Sue's age)

To solve this, we should make it so that there's only one variable to solve for.

Since Sue is 4 years older than David,
then let x be David's age and (x+4) be Sue's age.

Now, lets look at our facts.
22=x+y
x is David's age
y is Sue's age
x+4 is Sue's age

From here, we can conclude that y=x+4.
So from 22=x+y, we can substitute y for y=x+4.

22=x+y
22=x+(x+4)
22=2x+4

Now, solve for x.
22=2x+4
18=2x
9=x

Now, we have figured out David's age since x stands for David's age.
To solve for Sue's age, plug in 9=x for y=x+4

y=x+4
y=(9)+4
y=13

Now, we have found out Sue's age.
To recap, David is 9 years old and Sue is 13 years old.